# JEL classification: C33, C38, C55

### v1

@@
This paper proposes a novel method for sparse latent factor modeling using a new sparse asymptotic Principal Component Analysis (APCA). This approach analyzes the co-movements of large-dimensional panel data systems over time horizons within a general approximate factor model framework. Unlike existing sparse factor modeling approaches based on sparse PCA, which assume sparse loading matrices, our sparse APCA assumes that factor processes are sparse over the time horizon, while the corresponding loading matrices are not necessarily sparse. This development is motivated by the observation that the assumption of sparse loadings may not be appropriate for financial returns, where exposure to market factors is generally universal and non-sparse. We propose a truncated power method to estimate the first sparse factor process and a sequential deflation method for multi-factor cases. Additionally, we develop a data-driven approach to identify the sparsity of risk factors over the time horizon using a novel cross-sectional cross-validation method. Theoretically, we establish that our estimators are consistent under mild conditions. Monte Carlo simulations demonstrate that the proposed method performs well in finite samples. Empirically, we analyze daily stock returns for a balanced panel of S&P 500 stocks from January 2004 to December 2016. Through textual analysis, we examine specific events associated with the identified sparse factors that systematically influence the stock market. Our approach offers a new pathway for economists to study and understand the systematic risks of economic and financial systems over time.
Keywords: Factor Analysis, Asymptotic Principal Components, Sparsity, Large-Dimension, Panel Data
JEL classification: C33, C38, C55
@@ 1 Introduction
In the realm of big-data analysis, the study of large-dimensional panel data has gained significant prominence in recent years. Large-dimensional panel data refers to datasets where observations are recorded for multiple individuals or entities over time, resulting in a wealth of information over the spatial and time horizons. Analyzing such datasets is essential for understanding complex economic, financial, and social phenomena. However, as the dimensionality of the data increases, traditional statistical techniques encounter numerous challenges, including multicollinearity, computational complexity, and difficulties in extracting meaningful insights. To address these challenges, latent factor modeling, also referred to as statistical factor modeling in Campbell, Lo, and MacKinlay (1997), has emerged as a powerful and versatile approach in the field of panel data analysis. Latent factor models offer a structured framework for dimension reduction and capturing common sources of variation in high-dimensional data. For example, asset returns in finance are often modeled as functions of a small number of factors; see Ross (1976) and Connor and Korajczyk (1986, 1988). Macroeconomic variables of multiple countries are found to have common components; see Stock and Watson (1989), Gregory and Head (1999), and Forni, Hallin, Lippi, Reichlin, et al. (2000). In demand systems, the Engle curves can be expressed in terms of a finite number of factors; see Lewbel (1991). As the dimensions of the panel data systems increase, various factor models have been developed to reduce the dimensionalities under different scenarios. See, for example, the approximate factor models in Chamberlain and Rothschild (1983), Stock and Watson (2002b, a), Bai and Ng (2002), Bai (2003), Fan, Liao, and Mincheva (2013), Lettau and Pelger (2020), Pelger (2020), and Gao and Tsay (2023, 2024), and the dynamic factor models in Forni, Hallin, Lippi, and Reichlin (2000), among others. These models are particularly well-suited for large-dimensional panel datasets where the number of variables (or features) can be moderately large.
@@
One of the most profound challenges in factor modeling is to handle high-dimensional time series data efficiently and effectively. Sparse factor modeling represents a groundbreaking approach to this challenge, offering a powerful technique to extract meaningful insights from complex datasets characterized by a multitude of variables. Sparse factor modeling seeks to identify and capture the underlying structure of data while simultaneously promoting simplicity by selecting only a subset of relevant variables (or features) from the original set. The central idea behind sparse factor modeling is to uncover latent factors that drive the observed data’s variation while enforcing a degree of sparsity, meaning that only a few variables are deemed essential to explain the data’s structure. This approach is particularly valuable in scenarios where the number of variables far exceeds the number of observations, as it not only reduces computational demands but also enhances the interpretability of the results. Examples of articles concerning this approach include the sparse PCA in Jolliffe, Trendafilov, and Uddin (2003), Zou, Hastie, and Tibshirani (2006), Shen and Huang (2008), and Witten, Tibshirani, and Hastie (2009), and the sparse factor analysis with sparse loadings in Kristensen (2017), among others, where all of the aforementioned works are in line with factor analysis with sparse loadings, implying that each factor process is a linear combination of a small subset of the original panel series only.
@@
It is widely known that the key idea in a large-dimensional factor model is that the dimensions of both the timeline and the cross-section of the data are large, and most of the co-movements can be explained by a few factors. These factors and loadings are usually estimated by the conventional principal component analysis (PCA) method due to its ability to parsimoniously capture much of the information in a large number of variables. However, the resulting latent PCs or factors are typically linear combinations of of all cross-sectional units/variables, which are usually hard to interpret. The sparse PCA restricts the cardinality of the weight vectors for the PCs, or equivalently, the loadings for the factor processes, so the the PCs are sparse linear combinations of the underlying variables. Consequently, the PCs or factors are only linear combinations of a small subset of the cross-sectional units, which facilitates the interpretations of the resulting PCs or factors. Nevertheless, there is little literature concerning the justificaton of the sparsity assumption in the loadings. For example, the assumption may not be appropriate for financial returns, where the exposure of the returns to a market factor is universal and non-sparse, as pointed out in Pelger and Xiong (2022), see also the empirical evidence in Gao and Tsay (2021).
@@
In view of the above discussion, this paper marks a further development in the sparse factor modeling of large-dimensional panel data from a different perspective. Motivated by the empirical success of a general approximate factor model, see Pelger (2020) and the references therein, we also adopt such a similar framework. Unlike the sparse factor modeling approach with sparse loadings in existing literature, we assume the latent factor processes are sparse over time, while the loadings can be nonzero in general settings. Specifically, for a given time period of economic or financial series, we postulate that cross-sectional units may exhibit varying strengths of co-movement over time. During certain periods, the co-movement may not be significant enough, allowing idiosyncratic terms to dominate the co-movement, potentially resulting in the absence of a systematic risk factor. See the empirical evidence of the dominance of idiosyncratic risk in Campbell, Lettau, Malkiel, and Xu (2001) and Goyal and Santa-Clara (2003), among others. Consequently, the loading vector of each panel series can be treated as a linear combination of the corresponding series over the entire timeline. When a factor process is sparse, the corresponding loading associated with each series is a linear combination of the series over a small subset of the entire time horizon. This implies that only specific data points contribute to the strength of the loading for each series. The interpretation of this approach is quite straightforward. For example, asset returns in finance may not always exhibit systematic co-movements over the entire timeline, even though the market beta associated with the market factor can be significantly different from zero since it is calculated based on the entire timeline. However, clear and significant co-movement patterns do emerge during certain periods, such as when the Federal Reserve raises interest rates to combat inflation or during financial crises triggered by currency issues, debt, banking problems, or subprime mortgage defaults. Therefore, our proposed approach can facilitate researchers in economics and finance in discovering the connections between specific policies or events and the co-movements of economic or financial systems.
@@
In this paper, we first introduce approximate factor models and their estimation procedures using the conventional PCA or the asymptotic PCA (APCA) method introduced by Connor and Korajczyk (1986, 1988). We then present a sparse APCA technique by formulating a sparse factor modeling framework with sparse factors over time horizons. Under the $\ell_{0}$-constraint imposed on the factors, we propose a truncated power method to estimate the sparse factors in the one-factor case, corresponding to the hard-thresholding technique in the machine learning literature. Additionally, we develop a sequential deflation estimation procedure for multi-factor cases, utilizing the truncated power method along with certain projection techniques. The theoretical properties of the proposed estimators are established under some mild conditions when both the dimension and the sample size go to infinity. Simulated and real data examples are used to illustrate the proposed method. To the best of our knowledge, this is the first study in the literature to consider sparse factors over the timeline, and we establish theoretical guarantees for the estimation procedure. The proposed approach provides an effective way to bridge time and significant risk factors, which play a crucial role in studying the dynamics of large-dimensional panel systems.
In an empirical application, we apply the proposed method to study the dynamics of daily stock returns for a balanced panel of S&P 500 stocks from January 2004 to December 2016. Through textual analysis, we examine specific events associated with the identified sparse factors that systematically influence the stock market. Our findings reveal nine significant time factors that may systematically affect the financial market: Economic Indicators, Government Policies, Global Events, Market Sentiment, Company-specific Factors, Credit Risk, Oil Price, China, and Europe. The findings contribute to a deeper understanding of market dynamics and the critical events that drive stock market movements, providing a valuable resource for both investors and policymakers in navigating the complexities of financial markets. The empirical study suggests that our proposed method not only provides interpretable results but also establishes connections between co-movements and specific events over time, effectively bridging time and risk. Our approach offers a new pathway for economists to study and understand the systematic risks of economic and financial systems over time.
@@
Note that the proposed framework not only marks a clear distinction from existing sparse PCA or sparse factor modeling in the literature in terms of formulation and interpretation, but the technique used in this paper also differs from those employed in previous works. For example, in the sparse PCA literature, such as Zou, Hastie, and Tibshirani (2006), Shen and Huang (2008), and Witten, Tibshirani, and Hastie (2009), the true sparse eigenvectors are assumed to be constant vectors, and regularized estimation algorithms are proposed to estimate these sparse eigenvectors. In the realm of sparse factor modeling, such as in Kristensen (2017) and Uematsu and Yamagata (2022), the loadings are assumed to be sparse and non-random, with $\ell_{1}$-regularized methods used to estimate the sparse loadings. Conversely, in the sparse proximate factor modeling approach of Pelger and Xiong (2022), only the factor weights are assumed to be sparse, while the resulting loading and factor processes are non-sparse. In contrast with the aforementioned frameworks, the proposed sparse APCA method assumes that the random factor processes exhibit a sparse structure, meaning each factor data point can be either a random variable or zero over the timeline. A zero factor point indicates that the return series may not have co-movement at that specific time point. Our estimation procedure is developed under $\ell_{0}$-constraints on the factors. Therefore, our theoretical analysis differs from all the previously mentioned approaches, as the dependence structure along the timeline differs from that across the cross sections. Consequently, our method offers a way to link specific time points with the relevant risk factors that systematically influence the panel of interest.
This paper is also related to the literature on high-dimensional factor models with weak factors. In this context, there are two classes of weak factors: the first type affects only a subset of cross-sectional units, while the second type has relatively weak effects on the entire cross-section. The former type is often estimated using a group structure or some prior knowledge about factor strength, as discussed in Bailey, Kapetanios, and Pesaran (2021) and Anatolyev and Mikusheva (2022). In addition, Huang, Jiang, Li, Tong, and Zhou (2022) proposed a scaled PCA method to scale down weak factors through a supervised learning approach. This method was later extended to a supervised dynamic PCA method by Gao and Tsay (2024) for dynamic forecasting with weak factors. Importantly, Bai and Ng (2023) showed that PCA estimators exhibit a slower convergence rate under a weak factor model. Our paper fits into the second category, where factors over the time horizon are assumed to be either zero or nonzero, denoting weak and strong factors, respectively. Zero factors indicate that idiosyncratic components are relatively more dominant compared to weak systematic risk factors. For long time series sequences, our method also offers a dimension-reduction approach (in time horizon) to studying the dynamics of the economic and financial series over time horizons.
@@ 2.1 Model Setup
where $\boldsymbol{\Lambda}=(\boldsymbol{\lambda}_{1},...,\boldsymbol{\lambda}_{N})^{\prime}$ and ${\mathbf{e}}_{t}=(e_{1,t},...,e_{N,t})^{\prime}$. Assume a panel data set of $T$ time-series observations and $N$ cross-sectional observations, denoted as ${\mathbf{X}}\in\mathbb{R}^{T\times N}$, has a factor structure with $r$ common factors. Let ${\mathbf{F}}=({\mathbf{f}}_{1},...,{\mathbf{f}}_{T})^{\prime}$ and ${\mathbf{e}}=({\mathbf{e}}_{1},...,{\mathbf{e}}_{T})^{\prime}$, then Models (1) and (2) can be written in the following matrix form
@@
For ease of notation, we also denote ${\mathbf{X}}=(\underline{{\mathbf{x}}}_{1},...,\underline{{\mathbf{x}}}_{N})$, ${\mathbf{F}}=(\underline{{\mathbf{f}}}_{1},...,\underline{{\mathbf{f}}}_{r})$, and $\boldsymbol{\Lambda}=(\underline{\boldsymbol{\lambda}}_{1},...,\underline{\boldsymbol{\lambda}}_{r})$ when referring to their columns. In the econometrics/statistical and finance literature, Model (3) is usually estimated using the Principal Components or Asymptotic Principal Components estimation method (see Connor and Korajczyk (1986, 1988), Bai and Ng (2002), Bai (2003), and Fan, Liao, and Mincheva (2013), among others). As a result, all the elements in the estimators $\widehat{\mathbf{F}}$ and $\widehat{\boldsymbol{\Lambda}}$ are usually nonzero, making it difficult to interpret the dynamic and cross-sectional relationships of the data. Moreover, the role of the idiosyncratic terms is rarely considered or even ignored in characterizing the individual dynamics of the series.
@@ 2.2 Identification and PCA Estimation
$\textnormal{Cov}({\mathbf{f}}_{t})={\mathbf{I}}_{r}\,\,\text{and}\,\,\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}\,\,\text{is diagonal},$ (4)
@@
$\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}={\mathbf{I}}_{r}\,\,\text{and}\,\,\textnormal{Cov}({\mathbf{f}}_{t})\,\,\text{is diagonal}.$ (5)
@@
See, for example, Bai and Ng (2002) and Fan, Liao, and Mincheva (2013). Note that ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$ are still not identifiable when the diagonals in (4) or (5) have multiplicity or there exists a diagonal matrix ${\mathbf{H}}$ with the diagonals being $1$ or $-1$ such that $({\mathbf{F}}{\mathbf{H}},\boldsymbol{\Lambda}{\mathbf{H}})$ is still a pair of solutions. In this paper, we do not consider the special case of multiplicities, and the change of signs does not affect our analysis because it gives us an advantage to choose an ${\mathbf{F}}$ with any particular signs to facilitate the interpretations of the estimates. On the other hand, according to the discussion in Bai and Ng (2013), the estimates can be uniquely defined asymptotically under conditions (4) or (5) if the top eigenvalues of the covariance of the panel are distinct. They showed that the rotation matrix ${\mathbf{H}}$ is asymptotically equal to an identity with certain convergence rates. As a matter of fact, the conditions in (4) and (5) are reasonable if we adopt the PCA method to estimate the factors and their associated loadings, because the estimated PCA factors $\widehat{\mathbf{F}}$ and loading $\widehat{\boldsymbol{\Lambda}}$ either satisfy the condition in (4) or (5) depending on whether we apply PCA to ${\mathbf{X}}{\mathbf{X}}^{\prime}$ or ${\mathbf{X}}^{\prime}{\mathbf{X}}$ as discussed below.
Suppose ${\mathbf{x}}_{t}$ is a demeaned vector, under condition (4), the estimation of ${\mathbf{F}}$ is obtained by conducting a PCA on ${\mathbf{X}}{\mathbf{X}}^{\prime}$ or a scaled version $\frac{{\mathbf{X}}{\mathbf{X}}^{\prime}}{NT}$, and each estimated loading vector $\boldsymbol{\lambda}_{i}$ is a linear combination of the $i$th series $\underline{{\mathbf{x}}}_{i}=(x_{i,1},...,x_{i,T})^{\prime}$ using the weight matrix $\widehat{\mathbf{F}}$. Specifically, under condition (4), the estimated factor matrix, denoted by $\widehat{\mathbf{F}}$, is $\sqrt{T}$ times the normalized eigenvectors corresponding to the $r$ largest eigenvalues of the $T\times T$ matrix ${\mathbf{X}}{\mathbf{X}}^{\prime}$. For a given $\widehat{\mathbf{F}}$, the estimated factor loading matrix is $\widehat{\boldsymbol{\Lambda}}^{\prime}=\widehat{\mathbf{F}}^{\prime}{\mathbf{X}}/T$, which can be obtained by ordinary least-squares (OLS) estimation. This approach is also known as Asymptotic Principal Component Analysis (APCA) in the finance literature; see Connor and Korajczyk (1986, 1988) for details. Consequently, the estimated loading vector $\boldsymbol{\lambda}_{i}$ for the $i$-th series is a linear combination of the $i$-th series $\underline{{\mathbf{x}}}_{i}=(x_{i,1},...,x_{i,T})^{\prime}$ using the weight matrix $\widehat{\mathbf{F}}$ over the timeline.
Similarly, under condition (5), the estimation of $\boldsymbol{\Lambda}$ is obtained by performing a PCA on ${\mathbf{X}}^{\prime}{\mathbf{X}}$ (or $\frac{{\mathbf{X}}^{\prime}{\mathbf{X}}}{NT}$). By abuse of notation, we still denote the estimated loading and factor matrices as $\widehat{\boldsymbol{\Lambda}}$ and $\widehat{\mathbf{F}}$, respectively. The estimated loading matrix $\widehat{\boldsymbol{\Lambda}}$ is constructed as $\sqrt{N}$ times the normalized eigenvectors corresponding to the $r$ largest eigenvalues of the $N\times N$ matrix ${\mathbf{X}}^{\prime}{\mathbf{X}}$. Then, the estimated factor matrix is $\widehat{\mathbf{F}}={\mathbf{X}}\widehat{\boldsymbol{\Lambda}}/N$. Therefore, each factor $\widehat{\mathbf{f}}_{t}$ is a linear combination of ${\mathbf{x}}_{t}=(x_{1,t},...,x_{N,t})$ using the weights in $\widehat{\boldsymbol{\Lambda}}$ over the cross section.
We highlight that the conditions in (4) and (5) are only imposed in order to estimate the factors and loading matrix by the PCA method, and the true underlying factors and loading matrix do not necessarily satisfy either (4) or (5) so long as some regularity conditions as those in Assumptions A-B of Bai and Ng (2002) are met. For example, if the estimated factors satisfy that $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T={\mathbf{I}}_{r}$, then there exists an invertible rotation matrix ${\mathbf{H}}\in R^{r\times r}$ such that the consistency of $\widehat{\mathbf{F}}$ to ${\mathbf{F}}{\mathbf{H}}^{\prime}$ can be established; See Bai (2003) and Fan, Liao, and Mincheva (2013) for details. Moreover, if we further assume the true factors satisfy ${\mathbf{F}}^{\prime}{\mathbf{F}}/T={\mathbf{I}}_{r}$, or more rigorously, $\textnormal{Cov}({\mathbf{f}}_{t})={\mathbf{I}}_{r}$ as that in (4), then the rotation matrix ${\mathbf{H}}$ is an orthogonal one (asymptotically) as shown in Fan, Liao, and Mincheva (2013). If we additionally assume the loading matrix satisfy the condition in (4), then ${\mathbf{H}}$ is an identity asymptotically, as shown in Bai and Ng (2013).

### v2

@@
This paper introduces a novel sparse latent factor modeling framework using sparse asymptotic Principal Component Analysis (APCA) to analyze the co-movements of high-dimensional panel data over time. Unlike existing methods based on sparse PCA, which assume sparsity in the loading matrices, our approach posits sparsity in the factor processes while allowing non-sparse loadings. This is motivated by the fact that financial returns typically exhibit universal and non-sparse exposure to market factors. Unlike the commonly used $\ell_{1}$-relaxation in sparse PCA, the proposed sparse APCA employs a truncated power method to estimate the leading sparse factor and a sequential deflation method for multi-factor cases under $\ell_{0}$-constraints. Furthermore, we develop a data-driven approach to identify the sparsity of risk factors over the time horizon using a novel cross-sectional cross-validation method. We establish the consistency of our estimators under mild conditions as both the dimension $N$ and the sample size $T$ grow. Monte Carlo simulations demonstrate that the proposed method performs well in finite samples. Empirically, we apply our method to daily S&P 500 stock returns (2004–2016) and identify nine risk factors influencing the stock market.
Keywords: Asymptotic Principal Components, Factor Analysis, Power Method, Sparsity, High-Dimension
@@ 1 Introduction
In the realm of big-data analysis, the study of large-dimensional panel data has gained significant prominence in recent years. Large-dimensional panel data refers to datasets where observations are recorded for multiple individuals or entities over time, resulting in a wealth of information over the spatial and time horizons. Analyzing such datasets is essential for understanding complex economic, financial, and social phenomena. However, as the dimensionality of the data increases, traditional statistical techniques encounter numerous challenges, including multicollinearity, computational complexity, and difficulties in extracting meaningful insights. To address these challenges, latent factor modeling, also referred to as statistical factor modeling in Campbell et al. (1997), has emerged as a powerful and versatile approach in the field of panel data analysis. Latent factor models offer a structured framework for dimension reduction and capturing common sources of variation in high-dimensional data. For example, asset returns in finance are often modeled as functions of a small number of factors; see Ross (1976) and Connor and Korajczyk (1986, 1988). Macroeconomic variables of multiple countries are found to have common components; see Stock and Watson (1989), Gregory and Head (1999), and Forni et al. (2000b). In demand systems, the Engle curves can be expressed in terms of a finite number of factors; see Lewbel (1991). As the dimensions of the panel data systems increase, various factor models have been developed to reduce the dimensionalities under different scenarios. See, for example, the approximate factor models in Chamberlain and Rothschild (1983), Stock and Watson (2002b, a), Bai and Ng (2002), Bai (2003), Fan et al. (2013), Lettau and Pelger (2020), Pelger (2020), and Gao and Tsay (2023, 2024), and the dynamic factor models in Forni et al. (2000a), among others. These models are particularly well-suited for large-dimensional panel datasets where the number of variables (or features) can be moderately large.
@@
One of the most profound challenges in factor modeling is to handle high-dimensional time series data efficiently and effectively. Sparse factor modeling represents a groundbreaking approach to this challenge, offering a powerful technique to extract meaningful insights from complex datasets characterized by a multitude of variables. Sparse factor modeling seeks to identify and capture the underlying structure of data while simultaneously promoting simplicity by selecting only a subset of relevant variables (or features) from the original set. The central idea behind sparse factor modeling is to uncover latent factors that drive the observed data’s variation while enforcing a degree of sparsity, meaning that only a few variables are deemed essential to explain the data’s structure. This approach is particularly valuable in scenarios where the number of variables far exceeds the number of observations, as it not only reduces computational demands but also enhances the interpretability of the results. Examples of articles concerning this approach include the sparse PCA in Jolliffe et al. (2003), Zou et al. (2006), Shen and Huang (2008), and Witten et al. (2009), and the sparse factor analysis with sparse loadings in Kristensen (2017), among others, where all of the aforementioned works are in line with factor analysis with sparse loadings, implying that each factor process is a linear combination of a small subset of the original panel series only.
@@
It is widely known that the key idea in a large-dimensional factor model is that the dimensions of both the timeline and the cross-section of the data are large, and most of the co-movements can be explained by a few factors. These factors and loadings are usually estimated by the conventional principal component analysis (PCA) method due to its ability to parsimoniously capture much of the information in a large number of variables. However, the resulting latent PCs or factors are typically linear combinations of all cross-sectional units/variables, which are usually hard to interpret. The sparse PCA restricts the cardinality of the weight vectors for the PCs, or equivalently, the loadings for the factor processes, so the PCs are sparse linear combinations of the underlying variables. Consequently, the PCs or factors are only linear combinations of a small subset of the cross-sectional units, which facilitates the interpretations of the resulting PCs or factors. Nevertheless, there is little literature concerning the justificaton of the sparsity assumption in the loadings. For example, the assumption may not be appropriate for financial returns, where the exposure of the returns to a market factor is universal and non-sparse, as pointed out in Pelger and Xiong (2022), see also the empirical evidence in Gao and Tsay (2021).
@@
In view of the above discussion, this paper marks a further development in the sparse factor modeling of large-dimensional panel data from a different perspective. Motivated by the empirical success of a general approximate factor model, see Pelger (2020) and the references therein, we also adopt such a similar framework. Unlike the sparse factor modeling approach with sparse loadings in existing literature, we assume the latent factor processes are sparse over time, while the loadings can be nonzero in general settings. Specifically, for a given time period of economic or financial series, we postulate that cross-sectional units may exhibit varying strengths of co-movement over time. During certain periods, the co-movement may not be significant enough, allowing idiosyncratic terms to dominate the co-movement, potentially resulting in the absence of a systematic risk factor. See the empirical evidence of the dominance of idiosyncratic risk in Campbell et al. (2001) and Goyal and Santa-Clara (2003), among others. Consequently, the loading vector of each panel series can be treated as a linear combination of the corresponding series over the entire timeline. When a factor process is sparse, the corresponding loading associated with each series is a linear combination of the series over a small subset of the entire time horizon. This implies that only specific data points contribute to the strength of the loading for each series. The interpretation of this approach is quite straightforward. For instance, asset returns in finance may experience periods of high volatility while also exhibiting relatively stable patterns during certain timeframes. It is known that clear and significant co-movement patterns do emerge during certain periods, such as when the Federal Reserve raises interest rates to combat inflation or during financial crises triggered by currency issues, debt, banking problems, or subprime mortgage defaults. Therefore, our proposed approach can facilitate researchers in economics and finance in discovering the connections between specific policies or events and the co-movements of economic or financial systems.
@@
In this paper, we first introduce approximate factor models and their estimation procedures using the conventional PCA or the asymptotic PCA (APCA) method introduced by Connor and Korajczyk (1986, 1988). We then present a sparse APCA technique by formulating a sparse factor modeling framework with sparse factors over time horizons. Under the $\ell_{0}$-constraint imposed on the factors, we propose a truncated power method to estimate the sparse factors in the one-factor case, corresponding to the hard-thresholding technique in the machine learning literature. Additionally, we develop a sequential deflation estimation procedure for multi-factor cases, utilizing the truncated power method along with certain projection techniques. Furthermore, a data-driven approach is also proposed to identify the sparsity of risk factors over the time horizon using a novel cross-sectional cross-validation method. Theoretical properties of the proposed estimators are established under mild conditions as both the dimension and sample size tend to infinity. Simulated and real data examples are used to illustrate the proposed method. We empirically examine specific events associated with the identified sparse factors that systematically influence the stock market. Our findings reveal nine significant time factors that may systematically affect the financial market. To the best of our knowledge, this is the first study in the literature to consider sparse factors over the timeline, and we establish theoretical guarantees for the estimation procedure. The proposed approach provides an effective way to bridge time and significant risk factors, which play a crucial role in studying the dynamics of large-dimensional panel systems.
@@
Note that the proposed framework not only marks a clear distinction from existing sparse PCA or sparse factor modeling in the literature in terms of formulation and interpretation, but the technique used in this paper also differs from those employed in previous works. For example, in the sparse PCA literature, such as Zou et al. (2006), Shen and Huang (2008), and Witten et al. (2009), the true sparse eigenvectors are assumed to be constant vectors, and regularized estimation algorithms are proposed to estimate these sparse eigenvectors. In the realm of sparse factor modeling, such as in Kristensen (2017) and Uematsu and Yamagata (2022), the loadings are assumed to be sparse and non-random, with $\ell_{1}$-regularized methods used to estimate the sparse loadings. Conversely, in the sparse proximate factor modeling approach of Pelger and Xiong (2022), only the factor weights are assumed to be sparse, while the resulting loading and factor processes are non-sparse. In contrast with the aforementioned frameworks, the proposed sparse APCA method assumes that the random factor processes exhibit a sparse structure, meaning each factor data point can be either a random variable or zero over the timeline. A zero-factor point suggests that the return series exhibits minimal co-movement or that prices remain relatively stable at that particular time. Our estimation procedure is developed under $\ell_{0}$-constraints on the factors. Therefore, our theoretical analysis differs from all the previously mentioned approaches, as the dependence structure along the timeline differs from that across the cross sections. Consequently, our method offers a way to link specific time points with the relevant risk factors that systematically influence the panel of interest.
@@ 2.1 Model Setup
where $\boldsymbol{\Lambda}=(\boldsymbol{\lambda}_{1},...,\boldsymbol{\lambda}_{N})^{\prime}$ and ${\mathbf{e}}_{t}=(e_{1,t},...,e_{N,t})^{\prime}$. Assume a panel data set of $T$ time-series observations and $N$ cross-sectional observations, denoted as ${\mathbf{X}}\in R^{T\times N}$, has a factor structure with $r$ common factors. Let ${\mathbf{F}}=({\mathbf{f}}_{1},...,{\mathbf{f}}_{T})^{\prime}$ and ${\mathbf{e}}=({\mathbf{e}}_{1},...,{\mathbf{e}}_{T})^{\prime}$, then Models (1) and (2) can be written in the following matrix form
@@
For ease of notation, we also denote ${\mathbf{X}}=(\underline{{\mathbf{x}}}_{1},...,\underline{{\mathbf{x}}}_{N})$, ${\mathbf{F}}=(\underline{{\mathbf{f}}}_{1},...,\underline{{\mathbf{f}}}_{r})$, and $\boldsymbol{\Lambda}=(\underline{\boldsymbol{\lambda}}_{1},...,\underline{\boldsymbol{\lambda}}_{r})$ when referring to their columns. In the econometrics/statistical and finance literature, Model (3) is usually estimated using the Principal Components or Asymptotic Principal Components estimation method (see Connor and Korajczyk (1986, 1988), Bai and Ng (2002), Bai (2003), and Fan et al. (2013), among others). As a result, all the elements in the estimators $\widehat{\mathbf{F}}$ and $\widehat{\boldsymbol{\Lambda}}$ are usually nonzero, making it difficult to interpret the dynamic and cross-sectional relationships of the data. Moreover, the role of the idiosyncratic terms is rarely considered or even ignored in characterizing the individual dynamics of the series.
@@ 2.2 Identification and PCA Estimation
$\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}={\mathbf{I}}_{r}\,\,\text{and}\,\,\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}\,\,\text{is diagonal},$ (4)
@@
$\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}={\mathbf{I}}_{r}\,\,\text{and}\,\,\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}\,\,\text{is diagonal}.$ (5)
@@
See, for example, Bai and Ng (2002) and Fan et al. (2013). Note that ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$ are still not identifiable when the diagonals in (4) or (5) have multiplicity or there exists a diagonal matrix ${\mathbf{H}}$ with the diagonals being $1$ or $-1$ such that $({\mathbf{F}}{\mathbf{H}},\boldsymbol{\Lambda}{\mathbf{H}})$ is still a pair of solutions. In this paper, we do not consider the special case of multiplicities, and the change of signs does not affect our analysis because it gives us an advantage to choose an ${\mathbf{F}}$ with any particular signs to facilitate the interpretations of the estimates. On the other hand, according to the discussion in Bai and Ng (2013), the estimates can be uniquely defined asymptotically under conditions (4) or (5) if the top eigenvalues of the covariance of the panel are distinct. They showed that the rotation matrix ${\mathbf{H}}$ is asymptotically equal to an identity with certain convergence rates. The conditions in (4) and (5) are reasonable when applying the PCA method to estimate factors and loadings. Specifically, the estimated PCA factors $\widehat{\mathbf{F}}$ and loadings $\widehat{\boldsymbol{\Lambda}}$ asymptotically satisfy either (4) or (5), depending on whether PCA is performed on ${\mathbf{X}}{\mathbf{X}}^{\prime}$ (Asymptotic PCA) or ${\mathbf{X}}^{\prime}{\mathbf{X}}$ (PCA method). For further details, see Connor and Korajczyk (1986, 1988) and Chapter 9 of Tsay (2010).